OpenStudy (ttp):

The point Negative twelve comma sixteen. would not fall on the line for which of the following representations? Seven minus two y equals two x minus 1. two y minus twenty-eight equals negative four x plus twelve. Negative three y equals six x plus twenty-four. Ten y equals negative five x plus one hundred.

OpenStudy (solomonzelman):

\[(-12,16)\] write each of the choices out in numbers first.

OpenStudy (ttp):

Okay A. 7- 2y= 2x -1 b. 2y - 28 = -4x +12 c. -3y =6x + 24 d. 10y= -5x + 100

OpenStudy (solomonzelman):

I would translate each into slope intercept formula, it looks like \[y=mx+b\]

OpenStudy (ttp):

So can you help me with that?

OpenStudy (solomonzelman):

A. y= x -4 b. y = -2x +20 c. y =-2x -12 d. y= -1/2 x + 10 Agree?

OpenStudy (ttp):

Agree

OpenStudy (solomonzelman):

you have a point\[(-12,16)\]I'll show on choice A, you do the rest. Plug it in. -12 for x, and 16 for y. \[y= x -4\]\[16= -12 -4,-->16≠-16\] So A can't be the answer, do the rest of the choices like me, OK?